AGw and GCS Commissioning Observing on 07/08 September 2008.

Observers: D. Miller, A. Rakich and J. Hill in Tucson. A herd of Lucifer people on the mountain
Telescope Operator: A. Ceranski

Summary

1 hour of the night for Lucifer Engineering

1 hour trying to get the irs and IIF to talk properly. Looked like OSS was having trouble, but later in the night we think it is not and OSS problem but an irs-IIF problem

3.5 hours lost to weather

3.5 hours taking field aberration data. Made two x -15 arcsec moves of the secondary in Rx. Marked improvement in evenness of field aberrations. May need one more iteration Monday night.

Nighttime Activities

3:30 Opened for Lucifer engineering

5:00 tried to preset telescope from IRAF to test guiding and Act Opt loop. IIF, irs problem. After 30 minutes of restarting a few things then called Chris B. and he found OSS was in a funky state. Once it was restarted then presetTelescope worked.

6:00 PresetTelescope with IRAF with guiding, but no guide star in acquisition image. Set to TRACK (BS9183) and set ie, ca to put star on hotspot

6:20 PresetTelescope with GUIDE and was successful. By the time has acquired a WFSC image, clouds has drifted overhead

6:25 Closed chamber because radar image showed rain in the area.

9:00 Open to mostly clear skies. pointto BS9107, TRACK. set

9:02 on-axis collimate 08,09

9:04 off-axis #29 Doug angle = -359 (want 0 and 360) Good! 10

9:10 on-axis collimate 11 not sent (irs frozen)

9:14 on-axis 12 sent

9:20 on-axis collimate 13 not sent (irs frozen)

9:25 on-axis 14-19 sent

9:30 off-axis # 29 pointto frozen, Aaron stop and start OSS. GSC then start guiding. Seems OSS is the problem IssueTrak #1559

9:35 on-axis 21-28 sent with occasional irs freezes (changed from DEFAULT to M1)

9:53 GUIDE on # 29 wfsc 29-31

10:00 on-axis (pointto GUIDE froze) collimate 32-33

10:04 off-axis #91 wfsc 34-36

10:10 on-axis TRACK (some clouds, on-axis star is dimmer)

10:14 M2 tilt -5, wrong direction, removed this -5

10:16 M2 tip -15, ie changed from -120 to -105

10:18 on-axis GUIDE, irs freeze

10:21 on-axis GUIDE, ok collimate 37-38

10:25 off-axis #91 wfsc 39-41

10:29 on-axis GUIDE, PA=275 collimate 42-43

10:33 off-axis #29 wfsc 44-46

10:36 on-axis wfsc 47-48

10:43 off-axis #60 irs frozen (too close to on-axis, 74 mm)

10:44 off-axis #69 wfsc 49-51

10:48 on-axis GUIDE PA=187 collimate 52

10:50 off-axis #68 wfsc 53-55

10:55 on-axis GUIDE PA=172 collimate 56-58 (59-62 problems with irs, not sent)

11:15 BS9108 on-axis GUIDE wfsc 63-64

11:22 off-axis #21 wfsc 65 no collimate

11:26 on-axis GUIDE PA=95 wfsc 66 collimate, 67-68 fail, 69 collimate

11:34 off-axis #16 failed

11:35 off-axis #16 wfsc 70-72 (72 failed to get DD)

11:41 on-axis GUIDE PA=0 collimate 73-74

11:45 off-axis #18 wfsc 75 failed 76-78 OK (78 sausages)

11:51 on-axis GUIDE PA=30 collimate 79-80

11:55 tip -15, ie changed from -105 to -90

11:59 on-axis GUIDE PA=30 collimate 81

12:02 off-axis #18 failed, to dim

12:04 off-axis #18 wfsc 82 failed 83-85 OK

12:09 on-axis GUIDE PA=180 collimate 86 failed 87 OK

12:15 off-axis wfsc 88-90

12:18 on-axis GUIDE PA= 180 failed

12:20 on-axis GUIDE PA=180 collimate 91

12:22 off-axis #6 failed

12:23 off-axis #6 wfsc 92-94

12:28 off-axis #21 94 failed dd, 95-96 ok 97 failed DD, 98

12:34, 35, 36 on-axis guides failed. to bright

Done for tonight

Final IE=-90, CA=-60

-- DougMiller - 08 Sep 2008
Topic revision: r4 - 12 Sep 2008, JohnHill
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback